OpenStudy (anonymous):

Find several points (x, y)that satisfy the equation. x f(x) = 5^−x -1= 1= 2=

OpenStudy (anonymous):

5^(-(-1)) = 5^1 = 5 5^(-1) = 1/5 5^(-2) = 1/25

OpenStudy (anonymous):

(-1,5) (1, 1/5) (2,1/25)